Fermi National Laboratory
 

Accelerator Update for May 2, 2007 - May 4, 2007

 

Wednesday May 2

The day shift began with Operations monitoring the Recycler stash (325.6E10), with the Antiproton Source (Pbar) stacking (10.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

At 8:15 AM, Operations put the Tevatron (TeV) into shot setup.

At 8:33 AM, a Linac Klystron RF station (KRF5) tripped off. It reset.

At 8:48 AM, a Linac RF station (LRF3) tripped off for the first of two times in about 30 minutes. It reset.

At 9:08 AM, a MI RF station (MIRF14) tripped off. It reset.

At 9:41 AM, Operations received a LCW leak alarm for the MI-60 magnet system. An expert investigated.

At 9:49 AM, MiniBooNE tripped off due to bad beam positions. It reset.

Operations established store 5402 at 10:18 AM with an initial luminosity of 245.99E30.

At 10:34 AM, Operations reported that MI vacuum problems were holding off MI beam. Operators contacted experts. Three vacuum valves had closed. Operators opened the valves and beam resumed at 10:41 AM.

At 11:02 AM, LRF3 tripped off. It reset.

Operations established beam to the new Meson MTest experiment ( T963) at 11:43 AM. The Brookhaven National Laboratory (BNL) T963 experimenters will use the Meson Test beam facility to test a Muon Telescope Detector (MTD), and the performance of Time-of-Flight system (TOF) electronics, multi-gap resistive plate chamber (MRPC) strips, and two wire chambers.

At 12:33 PM, the Linac Klystron Debuncher tripped off. It reset.

MI experts began a study at 1:15 PM.

Recycler experts reported at 2:35 PM that they installed a new board for their damper system that should resolve some of their emittance problems.

A Linac expert began working on the LRF4 phase regulator at 3 PM. Over the next hour LRF4 tripped off several times. In response he replaced a phase amplifier and module. He reported that the auto adjust wasn’t working and wouldn’t compensate for phase variations; more work would be needed.

Operations reported at 4:10 PM that the Recycler had stashed antiprotons twice during the day shift.

At 5:18 PM, the MiniBooNE beam permit tripped off three times over the next hour due to bad beam positions. The experiment was off for about 10 minutes after each trip before it was reset and turned on.

The MTest experiment finished using the beam for the day at 6:05 PM.

At 6:59 PM, KRF5 tripped off. It reset.

At 7:15 PM, Operations reported a Recycler emittance blow-up. Operations contacted an expert who said the problem was probably due to the new damper board. Later, he came in and changed back to the old board.

At 7:40 PM, the Linac beam intensity dropped. An expert was here and he investigated.

At 7:59 PM, LRF3 tripped off. It reset.

The Linac expert reported at 8:15 PM that the Linac Buncher had a bad power supply, which he replaced.

At 11:06 PM, KRF4 tripped off. It reset.

Operations reported at 11:57 PM that the Recycler had stashed three times during the evening shift.

Thursday May 3

The midnight shift began with Operations monitoring store 5402 (66.76E30) and the stash (200.8E10), with Pbar stacking (29.9E10), and with NuMI and MiniBooNE taking beam.

At 12:15 AM, a TeV injection kicker (T:E48K) tripped on an interlock fault. This device wouldn’t be needed until the next shot, so its repair will wait until morning.

At 2:59 AM, LRF3 tripped off. It reset.

At 6:16 AM, KRF3 tripped off. It reset.

Operations established beam to the MTest experiment at 6:25 AM.

TeV experts began upgrading the TeV sectors A1 BLM system.

Operations terminated store 5402 at 8:01 AM.

Operations reported at 8:08 AM that the Recycler had stashed three times during the midnight shift.

At 8:14 AM, the NuMI beam permit tripped off due to a bad beam position. It reset.

TeV experts began a four-hour sextupole study at 8:18 AM.

At 10:43 AM, LRF3 tripped off. It reset.

At 10:53 AM, KRF5 tripped off. It reset.

The Neutron Therapy Facility (NTF) reported at 11:22 AM that they had completed treatment on their last patient for the day.

Operations loaded the TeV with a 12x0 wet squeeze at 11:37 AM.

TeV studies ended at 12:52 PM.

Operations prepared CDF and D0 for access at 12:54 PM.

At 1:01 PM, LRF3 tripped off. An expert reset it.

At 1:04 PM, LFR4 tripped off. It reset.

At 1:16 PM, KRF3 tripped off. It reset.

Linac experts began work at 1:30 PM on LRF4 and accessed the I- Source dome to inspect equipment.

CDF and D0 were out of access by 3:24 PM.

Operations put the TeV through a dry squeeze at 3:34 PM.

At 3:55 PM, Operations reported that KRF3 was off and experts were investigating.

Operations reported at 3:59 PM that the Recycler had stashed once during the day shift.

Linac experts reported at 4:12 PM that LRF4 was operational.

Linac experts completed their work on a LCW regulation problem for KRF3 at 4:28 PM. Operations resumed sending beam to Pbar, NuMI, MiniBooNE, and MTest.

Operations put the TeV into shot setup at 4:34 PM.

At 4:46 PM, KRF3 LCW regulation problems returned; beam was intermittent.

At 5:29 PM, Linac experts reported that the KRF3 problem was due to a flow meter not reading properly. Experts adjusted the system.

At 6:11 PM, Operations reported problems with the MI-40 BPMs. Operators called in an expert.

At 6:23 PM, MIRF14 tripped off for the first of two times in 18 minutes. It reset.

Operations established store 5409 at 8:43 PM with an initial luminosity of 164.91E30.

At 10:31 PM, Operations reported a TeV sector B17 separator spark. Store 5409 lost about 15E30 of luminosity. Operators investigated.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Friday May 4

The midnight shift began with Operations monitoring store 5409 (90.38E30) and the stash (152.2E10), with Pbar stacking (12.4E10), and with NuMI and MiniBooNE taking beam.

At 1:18 AM, LRF3 tripped off. It reset.

At 5:28 AM, KRF4 tripped off. It reset.

Operations established beam to MTest at 6:12 AM.

At 6:24 AM, a Meson power supply (MT3Q2) shut itself off and then its reversing switch changed polarity. An expert reported that the supply does this when it sees a reset. He will put the reversing switch in local so it doesn’t reverse its polarity on a reset.

At 7:05 AM, KRF3 tripped off. It reset.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/04/2007    email Fermilab